OpenStudy (anonymous):

How to prove that a series is convergent by using Cauchy Convergence Criterion for Series?

OpenStudy (amistre64):

define the criterion

OpenStudy (amistre64):

im thining its: for e > 0 and m,n > N |xm - xn| < e

OpenStudy (anonymous):

The series \(\sum_{n=1}^\infty a_n\) converges if and only if for every \(\epsilon > 0\) there exists \(N_\epsilon> N\) such that for every \(n > m > N_\epsilon\) we have \(|a_{m+1} + a_{m+2} + + a_n| < \epsilon\)

OpenStudy (amistre64):

Cauchy’s criterion. The sequence xn converges if and only if for every e > 0 there exists a natural number N such that |xn − xm| < e whenever n, m > N.

OpenStudy (amistre64):

im not sure im familiar with your definition of it

OpenStudy (anonymous):

For Cauchy Convergence Criterion, there are two versions: one is for sequence, and the other one is for series. Yours is for sequence, mine is for series.

OpenStudy (amistre64):

hmm, then that would explain the differences :)

OpenStudy (amistre64):

so, what is your series that you are trying to cauchy?

OpenStudy (anonymous):

Well, I am not sure if I can post my homework problem. Let me just make up a similar one... \(\sum_{n=1}^\infty\frac{1}{n^2-5}\)

OpenStudy (amistre64):

so now we want to make a partial sum, it looks like

OpenStudy (anonymous):

Yes.

OpenStudy (ikram002p):

will show u ltr (after 3,4 hours ), if u still need it hehehe its much easy than u think :O

OpenStudy (amistre64):

if you have some guidance, feel free to participate :)

ganeshie8 (ganeshie8):

yeah basically you're just proving the sequence of partial sums converges

OpenStudy (anonymous):

@ikram002p What is the main idea of your approach?

ganeshie8 (ganeshie8):

\[\text{sequence of partial sums converges} \iff \text{series converges}\]

ganeshie8 (ganeshie8):

comparison using p-series is more natural for proving 1/(n^2-5) series though

OpenStudy (anonymous):

Here's what I've got so far: \(|S_n-S_{n-1}| = a_n\) For n>m, we can use triangle inequality to obtain \(|S_n-S_m| = |S_n - S_{n-1} + S_{n-1} + S_{n-2}+ ... +S_{m-2}-S_{m-1}+S_{m-1} -S_m|\) \(\le |S_n - S_{n-1}| + |S_{n-1} + S_{n-2}|+... +|S_{m-2}-S_{m-1}|+|S_{m-1} -S_m|\) \(=|a_n|+|a_{n-1} + ... + |a_{m-2}|+|a_{m-1}|\) Then... lost...

OpenStudy (anonymous):

Apart from proving it like that, I am trying to show that \(a_n\) is a montone sequence, so I would have \(a_n\) = convergent => \(a_n\) = cauchy, then done. But it doesn't seem to be a monotone sequence due to the first term...

OpenStudy (ikram002p):

u chosed n as n-1 and n as n , lets if its work , u need to expand a bit to make useful of scoping , it is monotone but u need to follow steps to prove using cauchy , hmm saying monotone is not enough , lets walk through the Def , again ok ?

ganeshie8 (ganeshie8):

well you can always get rid of first few terms with their partial sum : \[\large \sum\limits_{n=1}^{\infty}\frac{1}{n^2-5} = S_m+ \sum\limits_{n=m+1}^{\infty}\frac{1}{n^2-5} \]

OpenStudy (ikram002p):

but we should note its monotone after n=2 hmm buts thats also work

OpenStudy (anonymous):

If it is monotone and bounded, then it is convergent. A convergent sequence is a Cauchy sequence. Then we can use it to show that its series is also convergent by the Cauchy Convergence Criterion for series. Is it right?

OpenStudy (anonymous):

The trickiest part is... in my lecture notes, we have to include the case when n=1 for a sequence to be a monotone sequence.

OpenStudy (ikram002p):

ok go like this , chose as u said m=n+1 then m>n |s_(n+1) - s_(n) |= |x_1+x_2 +( the rest go with scoping ) |

ganeshie8 (ganeshie8):

telescoping ?

OpenStudy (anonymous):

@ganeshie8 How and what to telescope?

ganeshie8 (ganeshie8):

no, im asking ikram if she has meant to use telescoping for the sum inside absolute bars

OpenStudy (ikram002p):

yeah hehe but scope means chicking , for our luck its telescope here xD

OpenStudy (ikram002p):

checking*

OpenStudy (anonymous):

What is x_1?

OpenStudy (anonymous):

very interesting

OpenStudy (ikram002p):

its the value of x_n when u apply 1 like x_1= 1/(1)^2-5=1/-4 x_2=1/(2)^2-5 + 1/-4 x_n=.....

OpenStudy (anonymous):

Then no, |s_(n+1) - s_(n) | is not equal to |x_1+x_2 +...|, but ||s_(n+1) - s_(n) | = x_(n+1)

OpenStudy (ikram002p):

hehe only in this case :3

OpenStudy (ikram002p):

u've already chosed m=n+1

OpenStudy (anonymous):

If m = n+1, then it violates my assumption that n>m

OpenStudy (ikram002p):

i said m>n :O

OpenStudy (ikram002p):

its ok its just a matter of ordering , i worked on m>n>N(epsilon ) u worked on n>m>N(epsilon ) i dint note that sry :3

OpenStudy (anonymous):

Can you please explain this part then? |s_(n+1) - s_(n) |= |x_1+x_2 +( the rest go with scoping ) |

OpenStudy (anonymous):

\(S_{n+1} = \sum_{i=1}^{n+1}x_i = x_1+x_2+...+x_n+x_{n+1}\) \(S_{n} = \sum_{i=1}^{n}x_i = x_1+x_2+...+x_n\) \(S_{n+1}-S_{n} = x_{n+1}\) This is what I get

OpenStudy (ikram002p):

oh wait , thats right xD

OpenStudy (anonymous):

And I don't see how much it differs from my working, despite that I have done more than that.

OpenStudy (ikram002p):

ugh i told u u worked right from the start just expand xD then 1/(n+1)^2-5 = 1/(n^2+2n-4) <2

OpenStudy (anonymous):

How do you get this: 1/(n^2+2n-4) <2?

OpenStudy (ikram002p):

its monotone right ? sup of -1 |-1|=1 take any sup value >1 and make it epsilon u could chose \(\epsilon_0 + 1\)

OpenStudy (ikram002p):

sup= supremum

OpenStudy (anonymous):

No, it is NOT monotone

OpenStudy (anonymous):

For monotone, it is either increasing or decreasing for ALL \(n\in \mathbb{N}\)

ganeshie8 (ganeshie8):

first few terms should not stop you from applying convergence tests its the tail that matters

OpenStudy (ikram002p):

not series only x_ n+1 -1, 1/4, .... xD -1 not in terms and also 2 > -1 xD

OpenStudy (anonymous):

-1/4, -1, 1/4 Not increasing nor decreasing

OpenStudy (anonymous):

I understand that the first few terms doesn't matter for the convergence of the whole sequence/series, but what bugs me is when I put it in the proof, it is not that persuasive at all, especially when you claim it is monotone while it is not.

OpenStudy (ikram002p):

i might made a typo x_n+1=1/(n^2+2n-4) =-1,1/4,1/11., ... how about now monotone or what O.O

OpenStudy (ikram002p):

you just x_n+1 @ganeshie8 , hehe anything better than cauchy , even thought its the fastest way to say its converge xD

OpenStudy (anonymous):

Perhaps not. Maybe the fastest way to say it converge is by the comparison test.

OpenStudy (ikram002p):

that also my fav :O but cachy is for prof :3

OpenStudy (anonymous):

x_n+1=1/(n^2+2n-4) =-1,1/4,1/11., ... ^ still not monotone

OpenStudy (ikram002p):

anyay say its monotone for n>1 , thus take any value >1 (since first term is 1) and make sure it goes will with other terms the idea is not about monotone or not , its about finding epsilon that make any any value of the sequence x_n+1 < epsilon

OpenStudy (ikram002p):

do u have a problem by saying for n>1 its monotone ?

OpenStudy (ikram002p):

when u start it from 1/4,..... (PS we are working on x_n+1 = 1/n^2+2n-5) )

OpenStudy (ikram002p):

take sup =1/4 for example or 1 let it be 1

OpenStudy (anonymous):

No, if I can say it is monotone for n>1 and ignore the case when n=1, I know how how to finish this problem. Just show it is monotone + bounded => convergence => Cauchy => done

OpenStudy (ikram002p):

ugh

OpenStudy (ikram002p):

u said u wanna CC for series

OpenStudy (anonymous):

Yes, but CC for series came from CC for sequence

OpenStudy (ikram002p):

if u wanna show its monotone + bounded its a lema like in other question , if its been telling u show its cauchy then u can say its bounded + monotone done

OpenStudy (ikram002p):

haha i dont know if u see a difference xD but u will in first exam^_^ be carful :P

ganeshie8 (ganeshie8):

but RolyPoly is particular about the first two terms

ganeshie8 (ganeshie8):

he/she wants to use the sequence from 1->infinity

OpenStudy (ikram002p):

so ?

ganeshie8 (ganeshie8):

so the given sequence as-it-is is not monotone

ganeshie8 (ganeshie8):

because it has few bad terms in the start

ganeshie8 (ganeshie8):

lol im fine skipping past first few terms if my only goal is to test convergence

ganeshie8 (ganeshie8):

but if you want to write a neat proof, you will need to think all the corners

OpenStudy (anonymous):

^Exactly

OpenStudy (ikram002p):

yes

OpenStudy (ikram002p):

however its ur choice at the end :)

ganeshie8 (ganeshie8):

im getting late for dinner, brb

OpenStudy (anonymous):

Are there no ways to fix this problem though?

ganeshie8 (ganeshie8):

leme tag @nerdguy2535 before i leave

OpenStudy (ikram002p):

my way is like this :- 1- reduce the series into limited sequences i like m=n+1 ,m>n>=N(epsilon ) |s_m-s_n|= x_(n+1) ( all ur work ) 2- according to CC for series for all epsilon >0 , there exist N(epsilon) in n such that if m>n>=N(epsilon), then |s_m-s_n| =|...| <epsilon so , |x_(n+1)|= |1/n^2+2n-4| =|1,1/4,1/11,.... 1/n^2+2n-4| < epsilon take \( \epsilon = \epsilon_0 + 1 , \epsilon_0>1 \) hehehe xD @nerdguy2535 help lol

OpenStudy (anonymous):

Whats the exact question? Trying to show that the series: $$\sum_{n=1}^\infty\frac{1}{n^2-5}$$converges using the Cauchy Criterion?

OpenStudy (ikram002p):

yeahh

OpenStudy (anonymous):

I think I saw this somewhere earlier in the discussion. If \(n>m\), and \(S_n=\sum_{k=1}^n \frac{1}{k^2-5}\)then: $$\left|S_n-S_m\right|=\left|a_{m+1}+a_{m+2}+\cdots +a_n\right|$$$$\le |a_{m+1}|+|a_{m+2}|+\cdots +|a_n|$$using the triangle inequality. Is everyone up to here?

OpenStudy (ikram002p):

waiting u hehe

ganeshie8 (ganeshie8):

yes

OpenStudy (ikram002p):

and Joseph went off xD

OpenStudy (anonymous):

(i'm kinda in a hurry so I dont have too much time unfortunately). The idea is that since \(k^2-5<(k+1)^2-5\) for all k, we have: $$\frac{1}{(k+1)^2-5}<\frac{1}{k^2-5}.$$ So: $$|a_{m+1}|+\cdots +|a_n|\le (n-m)\cdot |a_{m+1}|.$$

OpenStudy (anonymous):

Try and show that as m goes to infinity, that $$(n-m)\cdot |a_{m+1}|\rightarrow 0$$

OpenStudy (anonymous):

Call n-m=k for some fixed k (since n>m).

OpenStudy (anonymous):

So you have: $$\left|\frac{k}{(m+1)^2-5}\right|,$$which you want to show goes to zero as m goes to infinity.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (ikram002p):

cool :O

OpenStudy (anonymous):

By forcing n>m, it suffices to just take m to go to infinite, because it will also take n to infinity. n=m+k after all.

OpenStudy (experimentx):

\[ \left| \frac{1}{n^2-5} + \dots + \frac{1}{m^2-5} \right | < \left| \frac{1}{n^2} + \dots + \frac{1}{m^2} \right | \le \left| \frac 1 {n^2} + \frac{1}{n(n+1)} + \dots +\frac {1}{m(m-1)}\right| \\ = \left| \frac 1 {n^2} + \frac 1 n - \frac {1}{m}\right| < \frac{1}{n^2} + \frac 1 n = \epsilon \]

OpenStudy (anonymous):

Ah! I'll try to work on it later. It's too late in my place and I have to go. Sorry!